OpenStudy (alprincenofl):

what is the decimal expansion of the integer that ( 101011111)2 as its binary expansion .

OpenStudy (alprincenofl):

what is the decimal expansion of the integer that \[( 101011111)_{2}\] as its binary expansion .

OpenStudy (anonymous):

\[ (101011111)_2=\\1\times 2^7+0\times 2^6+1\times 2^5+0\times 2^4+1\times 2^3+1\times 2^2+1\times 2^1+1\times 2^0 \]

OpenStudy (anonymous):

\[ =128 + 32+8+4+2+1 \]

OpenStudy (anonymous):

\[=(351)_{10}=3\times10^2+5\times10^1+1\times10^0\]

OpenStudy (alprincenofl):

\[2^{8} + 2^{6} + 2^{4} + 2^{3} + 2 + 1 = 351\]